OpenStudy (anonymous):

1/cosecA-cotA-1/sinA=1/sinA-1/cosec+cotA

OpenStudy (anonymous):

?

OpenStudy (anonymous):

one side is negative the other side, whats your point

OpenStudy (anonymous):

prove it

OpenStudy (anonymous):

cnt gtng it!!!!!!

OpenStudy (dumbcow):

combine fractions \[\frac{2}{\sin x} = \frac{2\csc x}{(\csc x + \cot x)(\csc x - \cot x)}\] cross-multiply, 2's cancel \[\csc^{2} x - \cot^{2} x = \sin x *\csc x = 1 \] \[\csc^{2} x - \cot^{2} x = \frac{1}{\sin^{2} x} -\frac{1}{\tan^{2} x} = \frac{1-\cos^{2} x}{\sin^{2} x} = 1\] thus proving equivalence